OpenStudy (anonymous):

Which of the following formulas may be used to generate the sequence, 12, 19, 26, 33, 40, …?

OpenStudy (dan815):

change is 7, tarting at 12

OpenStudy (dan815):

y=mx+b m=7 b=12<--- when x =0, b must be 12 y=7x+12, note x=0, gives your first term, if u want x=1 to be first term then Y=7(x-1)+12